OpenStudy (anonymous):

you split $1500 between two savings accounts.Account A pays a annual of 5% interest and account B pays for 4% annual interest. After one year, you have earned a total of $69.50 in interest. How much money did you invest in each account.

OpenStudy (anonymous):

help

OpenStudy (tkhunny):

Please NAME the amounts assigned to each account.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

A + B = 1500 0.05 A + 0.04 B = $69.50 Solve by first temporarily eliminating A or B, so multiply the bottom equation by 20, then subtract it from the top. That will give you an equation with B not A, which you can solve. Lastly, the first equation will then give you A from A = 1500 - B.

OpenStudy (anonymous):

ok thnaks

OpenStudy (tkhunny):

Seriously, NAME the amounts. dwc did not do this. I guess we are to assume that we know what "A" and "B" are. Please learn to WRITE clear and concise definitions.

OpenStudy (anonymous):

thanks i meant to ask why do you have to multiply by 20

OpenStudy (anonymous):

I thought the accounts were named: "A" and "B."

OpenStudy (anonymous):

i didny get any accounts i just gave you the question i got for my homework that is the exact question i got

OpenStudy (anonymous):

I chose to multiply by 20 to get the coefficient in front of A in the second equation = 1, so I could then subtract the two equations and get rid of A temporarily.

OpenStudy (anonymous):

0.05 is turned into 1 you get a

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Once 0.05 is made into 1 etc, you can subtract the new equation from the top one and get B alone.

OpenStudy (anonymous):

then you will be left with 0.08 = 69.50 but then what do i do

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

No, I'll work it here: A + B = 1500 0.05 A + 0.04 B = 69.50 1 A + 0.80 B = 69.50 x 20 = 1390 subtract 0.2 B = 1500 - 1390 = 110 so B = 5 x 110 = 550 A= 1500 - B = 950 check 0.05(950) + 0.04(550) = ? = 69.50 checks

OpenStudy (anonymous):

ohhh ok thanks b]